### v2

# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

Abstract
Language Models (LMs) continue to advance, improving response quality and coherence. Given Internet-scale training datasets, LMs have likely encountered much of what users may ask them to generate in some form during their training. A plethora of evaluation benchmarks have been constructed to assess model quality, response appropriateness, and reasoning capabilities. However, the human effort required for benchmark construction is rapidly being outpaced by the size and scope of the models under evaluation. Having humans build a benchmark for every possible domain of interest is impractical. Therefore, we propose a methodology for automating the construction of fact-based synthetic data model evaluations grounded in document populations. This work leverages the same LMs to evaluate domain-specific knowledge automatically, using only grounding documents (e.g., a textbook) as input. This synthetic data benchmarking approach corresponds well with human curated questions producing a Spearman ranking correlation of $0.97$ and a benchmark evaluation Pearson accuracy correlation of $0.75$. This novel approach supports generating both multiple choice and open-ended synthetic data questions to gain diagnostic insight of LM capability. We apply this methodology to evaluate model performance on two recent arXiv preprints, discovering a surprisingly strong performance from Gemma-3 models on open-ended questions. Code is available at https://github.com/mmajurski/grounded-synth-lm-benchmark
1 Introduction
Figure 1: Synthetic evaluation pipeline: (1) users curate grounding documents, (2) LMs generate domain-specific questions, (3) model responses are evaluated.
Recent advances in language models (LMs) have demonstrated significant capability improvements in diverse natural language processing tasks. This expansion in capability has primarily been enabled by training on Internet-scale corpora. As AI tooling improves professional workflows, rigorous evaluation is imperative to understand the domain-specific model capability.
Traditional multiple choice question (MCQ) benchmarks such as MMLU [1] offer valuable general insights into model performance, but are static, public, and limited in scope. The fact that these benchmarks are static and public—and therefore can appear in training data—is a significant weakness. Combined with the human effort required to construct new high quality tests [2], it is unlikely evaluation construction can keep pace with LM development. This scaling challenge is especially acute for domain-specific professional applications, where the general benchmarks provide limited insight. Unfortunately, creating expert-validated benchmarks for every domain is impractical. New automated methods must generate accurate, comprehensive, diagnostically relevant evaluations grounded in authoritative corpora with minimal human intervention beyond document selection.
We propose an novel automated methodology for generating synthetic benchmarks grounded in authoritative documents. This approach reduces human effort and enables rapid adaptation to new domains by generating fact-based multiple choice or open-ended questions directly from appropriate grounding documents. This work is organized around three core questions:
  1. 1.
    How can LMs extract relevant information and generate high-quality questions grounded in domain documents?
  2. 2.
    Do these synthetic benchmarks replicate existing human-curated evaluations?
  3. 3.
    What limitations require mitigation for successful synthetic data evaluation of LM capability?
2 Related works
Our approach is motivated by task asymmetry similar to NP-completeness: given grounding documents, an LM should reliably generate diagnostically relevant questions, while zero-shot answering the same questions is substantially harder. This information asymmetry motivates using LMs to generate benchmarks grounded in curated documents rather than relying only on model weight knowledge. Language models are routinely adopted as judges: extracting answer choices [3, 4], grading open ended responses [5, 6], and justifying correctness [7]. The most similar contemporary convergent work is YourBench [8] synthetic replication of MMLU from grounding documents and AutoBencher [9] reinforcement learning optimization to discover gaps in LM knowledge. Leveraging grounding documents greatly improves synthetic data utility, as the methodology no longer relies on internalized LM knowledge.
Modern long-context LMs can process entire documents. Benchmarks like SecQA [10] and PubMedQA [11] leverage domain-specific textbooks and web pages. Full-document grounding generally outperforms summary-based systems [12]. Continued LM improvement, prompting strategies [13], and topic breadth [14] has further improved synthetic data. These methods also enhance downstream applications like retrieval-augmented generation [15].
LMs can serve as acceptable feature extractors when guided by detailed prompts. Studies show strong correlation between LM-based and human judgments of text quality [16, 17]. This aligns with the LM-as-a-Judge paradigm used for open-ended benchmark evaluation [18, 19]. Beyond open ended answer correctness, LMs can extract other textual features. However, limitations in LM ground truthing remain, especially in tasks like information retrieval where relevance judgments are more complex [20, 21]. LM evaluation benchmarks generally use either multiple choice (MC) or open-ended (OE) formats. Open-ended questions require synthesis [22] but depend on a judge LM for scoring [18, 23], while MCQs are easier to grade but may inflate performance due to answer visibility [24, 25]. Frameworks like Inspect-AI [26] support both types with prompt templates and answer shuffling. Both formats are sensitive to rewording, leading to inconsistent outputs [27], and newer frameworks aim to improve open-ended scoring precision [28, 29].
3 Methods
Constructing synthetic data benchmarks from relevant domain-specific documents requires chunking into contexts, topic extraction per context, question and answer generation, and finally an evaluation framework to determine model capability on the newly minted benchmark. All LM interactions use default generation parameters, except we set $\text{temperature}=0.7$ (o4-mini did not support setting temperature).
Document chunking into contexts
While modern LMs have sufficient context windows to support whole document ingestion, chunking encourages the generated questions to be spread throughout the grounding document. An input document is first converted into markdown for interoperability with LMs. We use Docling [30] for conversion from pdf to markdown, and pandoc [31] for conversion from html to markdown. Once in markdown, each document section is used as a context chunk for question generation. This chunking method is simple and relies on the document author crafting relevant sections. Enhancements to chunking and document data preparation will likely produce overall system improvements, but were not required to demonstrate baseline performance in this study. For instance, our approach only leverages the textual data; formatting and including richer sources of information from table, charts, and images may improve evaluation quality.
Topic extraction
For each extracted document context chunk, an LM is used to extract and propose any topics present which make high quality difficult questions for evaluating domain expert knowledge about the information in context chunks. This topic extraction leverages the same LM as the question generation process. We explored using an embedding model to cluster topics and de-duplicate, but empirically found no satisfying improvements, as the topic generation by LM does a reasonably good job of de-duplication. From manual observations, the only duplicate topics stem from different source document context chunks. This duplication is acceptable as the question generation is grounded in the context chunk; any duplicate topics across context chunks will produce different questions.
Question generation
Given (a) the document context chunk and (b) a single topic for a question, the generating LM is prompted to construct a high quality and difficult evaluation question to probe the knowledge and understanding of an expert in the domain. During question generation, the LM is required to write the question, correct answer, and a justification/explanation for why the correct answer is in fact right. The answer justifications are not used in the evaluation benchmark and simply exist to guard against generating incorrect answers. In summary, the single grounding document is broken into chunks of context, per-context relevant topics are extracted, then per-context/topic pair, and a single evaluation question is created. This results in a list of questions with correct answers grounded in (and derived from) the source document. This example drawn from Squadv2 [32] shows a question generated by Llama-3.3-70B-Instruct grounded in the context:
1{"context": "In 1756 and 1757 the French captured forts Oswego and William Henry from the British. The latter victory was marred when Frances native allies broke the terms of capitulation and attacked the retreating British column, which was under French guard, slaughtering and scalping soldiers and taking captive many men, women and children while the French refused to protect their captives. French naval deployments in 1757 also successfully defended the key Fortress of Louisbourg on Cape Breton Island, securing the seaward approaches to Quebec.",
2
3"question": "In what year did French naval deployments successfully defend the key Fortress of Louisbourg on Cape Breton Island, securing the seaward approaches to Quebec?",
4
5"choices":{"A": "1755",
6 "B": "1756",
7 "C": "1757",
8 "D": "1758"},
9
10"answer": "C"}
Multiple choice vs open-ended
In this study, we explored both multiple choice and open-ended question generation, each requiring slightly differing LM prompts. While simply using the correct multiple choice answer option in an open-ended configuration works, we observed that open-ended questions enable more variety from the generating model if prompts are modified. One study goal was to determine if any differences manifested between multiple choice and open-ended questions for synthetic data benchmarking. The difference in response grading requirements between open and multiple-choice favors MCQs for fast and simple exploration of model weaknesses.
3.1 Benchmark evaluation
Given a list of questions and answers, an LM evaluation framework (Inspect-AI [26]) is used to determine performance for the models under evaluation. The benchmark consists of many questions with associated answers. The resulting benchmark evaluation accuracy is the average number of correct answers divided by the total number of questions. LMs are evaluated against both the synthetic evaluation benchmark and the original questions. The overall accuracy per evaluated model can be compared for both the synthetic and original questions across multiple datasets of grounding documents. The text of the various benchmark questions and answers is never directly compared, only the resulting LM performance. Figure 2 showcases the performance of synthetic multiple choice questions averaged across datasets and generating models, with a Spearman ranking correlation of $0.968$, demonstrating slight model over-performance on the synthetic benchmark.
Figure 2: Comparison of average LM performance across six datasets with the multiple choice question synthetic data benchmarks created by the ensemble of bolded model in Table 2.
4 Evaluation methodology
Evaluation of the quality and utility of the newly minted synthetic data benchmark built from grounding documents is done by comparing the synthetic data pipeline on datasets where human annotated questions and correct answers exist for known context paragraphs. This provides a directly comparable result, both for the overall benchmark accuracy (between human and grounded synthetic data) and for the rank ordering of those models that the benchmark is evaluated on. We leverage six Natural Language Processing (NLP) datasets where each entry contains a context paragraph, human written question, and correct answer. For each dataset a variety of LMs are used to generate the synthetic benchmark. Per-question quality features and benchmark-level question diversity is extracted to compare against the human written questions. Finally, an automatic quality control pass is completed to grade both question and answer correctness/validity, in order to investigate and characterize which benchmark-generating models can write valid question and answer pairs.
4.1 Datasets
Evaluating the validity of generated synthetic questions requires source documents which have human annotated question/answer pairs to compare against. The following six NLP datasets were selected: (a) Squadv2 [32] (b) HotpotQA [33] (c) TrivaQA-web [34] (d) NaturalQuestionsShort [35] (e) PubMedQA [11], and (f) SecQA [10].
Table 1: Example human question reformatting
Context: Super Bowl 50 was an American football game to determine the champion of the National Football League (NFL) for the 2015 season. The American Football Conference (AFC) champion Denver Broncos defeated the National Football Conference (NFC) champion Carolina Panthers 24–10 to earn their third Super Bowl title. The game was played on February 7, 2016, at Levi’s Stadium in Santa Clara, California. As this was the 50th Super Bowl, the league emphasized the “golden anniversary” and temporarily suspended Roman numerals in the logo.
Original Question Original Answer Reformated Question Answer Choices
What day was the game played on? February 7, 2016 On which specific date did Super Bowl 50 take place at Levi’s Stadium in Santa Clara, California? (A) January 7, 2016 (B) February 7, 2016 (C) February 14, 2016 (D) January 14, 2016
If Roman numerals were used, what would Super Bowl 50 have been called? Super Bowl L What would the 50th Super Bowl have been called under the traditional Roman-numeral naming? (A) Super Bowl XL (B) Super Bowl XLVIII (C) Super Bowl XLIX (D) Super Bowl L
These datasets collectively contain tens of thousands of context, question, answer tuples. A random subset of 100 tuples per dataset was sampled as the grounding context document. For certain datasets, pre-processing was required to clarify, curate, and reformat the source human written questions into self-contained questions. For some questions this required disambiguation so the question is answerable without the context. For the multiple choice evaluation, three believable incorrect answer options needed to be created. The PubMedQA [11] dataset required conversion from yes/no/maybe questions into multiple-choice or open-ended. Only the SecQA [10] dataset already contained all of the required information and was not pre-processed. All other datasets went through an LM based reformatting operation correct these deficiencies. Table 1 provides an example of the reformatting operation for a few questions drawn from a single context of the Squadv2 dataset.
Three different models (Llama-4-Maverick-17B-128E-Instruct, gemma-3-27b-it, o4-mini) were used for this reformatting operation. The remainder of this paper shows results based on the Llama-4-Maverick-17B-128E-Instruct reformatting (see Table 4 in Appendix E for results of all three reformatting models).
4.2 Question generation models
Table 2: Question generating language models
Google gemma-3-4b-it [36]
gemma-3-12b-it [36]
gemma-3-27b-it [36]
Meta Llama-3.2-3B-Instruct [37]
Llama-3.1-8B-Instruct [38]
Llama-3.3-70B-Instruct [38]
Llama-4-Maverick-17B-128E-Instruct [39]
OpenAI o4-mini [40]
gpt-4.1 [40]
gpt-4.1-mini [40]
Microsoft phi-4 [41]
Qwen Qwen2.5-7B-Instruct [42]
To explore the impact of model capability on successful generation of synthetic data benchmarks we tested a variety of language models during the synthetic benchmark generation phase (see Table 2). All open-weight models are available through HuggingFace [43] and the closed-weight models were accessed through paid APIs. Additional results are reported for an ensemble of the bolded models. During benchmark creation. the same model is used both for topic extraction and question generation.
4.3 Evaluating against human questions
The reformatted human annotated question/answer pairs are used to compare the grounded synthetic data benchmark performance against the human. Ideally, the synthetic benchmark from grounding documents would produce the same LM capability measurements as the human-curated benchmark for any given model under evaluation. As such, to validate our approach, various LMs of differing sizes are evaluated on both the human and synthetic benchmarks. Comparison of pure benchmark evaluation accuracy is performed using the Pearson accuracy correlation between human and synthetic benchmark accuracy numbers. Evaluating the comparative rank ordering of models uses the Spearman rank correlation of relative model performance on the given benchmark dataset. These two measures compare both absolute performance alignment and relative rank order preservation/stability.
4.4 Generated question quality characterization
Per-question quality features and benchmark dataset-level features are computed to determine the overall quality of the synthetic benchmark. Owing to the low level of human oversight and annotation time for creating the synthetic data benchmarking, LMs extract these somewhat imprecise features. This relies on the idea that with enough guidance LMs can be good alternatives to human evaluation [16, 17] and borrows from standard practice LLM-as-a-Judge grading of open-ended benchmark responses [18, 19]. Llama-4-Maverick-17B-128E-Instruct extracts the following information about each generated question:
  1. 1.
    question token count: number of tokens in the question
  2. 2.
    question clarity: how clear and comprehensible is the question in isolation without the associated context
  3. 3.
    question groundedness: how grounded is the question in the context
  4. 4.
    answer correctness: how correct is the answer given the context
  5. 5.
    answer explanation validity: how valid and logically consistent the answer explanation is
In addition to per-question measurements, question diversity is measured over each benchmark as the average cosine similarity between all questions (using sentence-transformers/all-MiniLM-L6-v2 as the embedding model). The answer_correctness and answer_explanation_validity numbers flag which generating models are incapable of producing diagnostically useful questions.
5 Results
All experiments and development were conducted on a single 8×H100 on-premise machine. Development consumed approximately twice the compute of final evaluations. Final compute was split between synthetic dataset generation and benchmark evaluation via Inspect-AI. The most expensive model, gpt-4.1, was only used during final results, processing 3.6M input and 1.1M output tokens ( \$16), with compute evenly divided between dataset creation and evaluation.
5.1 Answer correctness
Synthetic benchmark answers must be accurate, a task made easier via the grounding documents that supply the necessary information. To verify generated answer accuracy, an LM scored each question-answer pair and its justification (1-10 scale) using the full context. As shown in Figure 3, reasonably sized models scored above 9.5 in accuracy, with smaller models performing noticeably worse. While model size/capability generally correlates with desirable properties for synthetic data generation, notable exceptions (like gpt-4.1) are discussed later.
Figure 3: Synthetic data generated question correctness and explanation validity on a scale of 1 to 10.
5.2 Question clarity and groundedness
Figure 4: Generated question clarity and groundedness compared against the original questions.
Question clarity is inherently tied to difficulty, as simple questions have the potential to be clearer, easier to understand, and easier to answer. Conversely, a needlessly opaque question is harder to answer while providing minimal additional diagnostic relevance, though a model’s ability to reason through more opaque questions indicates general capability. This study found that most generated questions have acceptable clarity without significant variation between the generating models. Most generated synthetic questions are longer and more complicated than the original questions. This increased average question complexity combined with similar levels of question clarity suggests most synthetic questions are well written. Manual random spot inspection supports this as we observed few failures.
How grounded in the source material (context chunk) a given question is indicates how much outside information both the generation and evaluation models must leverage to correctly interpret what is being asked and answer. Low grounding indicates the question/answer pair cannot be found within the context chunk, or that significant external information is required. High grounding means all necessary information is found within the context chunk. Therefore the generated question is testing for information aligned with and present in the context. Figure 4 highlights the question clarity and groundedness scores for the question generating models. While most LMs generate similar clarity values, the trends in groundedness score are more complex. Some capable models (like Llama-4-Maverick-17B-128E-Instruct) score well, other capable models (like o4-mini) score worse, with smaller models in between (gemma-3-4b-it). We observed two grounding failure modes: either the model hallucinates irrelevant content absent from the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). See Appendix F for an example from phi-4. Models of medium size perform well on groundedness, simply reformatting information presented in the context.
5.3 Impact of question length
[…]
(a) Question token count by generating model. Most models are within $20\%$ of human, but some produce overly long questions.
(b) Scatterplot documenting evaluation accuracy over-performance on long questions. Models like gpt-4.1-mini plot well above the trend line.
Figure 5: Impact of question length on evaluation accuracy over-performance.
5.4 Benchmark question diversity
During synthetic question generation, the explicit topic extraction step provides a backstop against degenerate questions which only focus on the most obvious elements of a context chunk. To compare the variety and coverage of the generated questions we compute the average cosine similarity between all question pairs (excluding self pair) over the generated synthetic dataset. While most generating models produce similar diversity scores closely aligned with the human questions, some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 in Appendix C shows the question diversity per generating model along with the original question diversity.
5.5 Benchmark model accuracy correlation
Figure 6: Benchmark evaluation accuracy correlation generated by model ensemble demonstrating reasonable correlation between human and synthetic evaluation. Pearson accuracy correlation $0.745$.
Ideally, we want synthetic benchmarks to correlate well with the evaluation accuracy measured by human curated questions. This ends up being true for some generating models, but not all. Figure 6 shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the five bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of $0.745$ between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 in Appendix E showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near $100\%$ evaluation accuracy (Pearson accuracy correlation $0.402$).
5.6 Benchmark model rank correlation
While direct Pearson correlation between human and synthetic benchmark evaluation accuracy is desirable, evaluations are often used to compare models. Therefore, a more important metric is the model rank order stability measured by Spearman rank correlation. Figure 2 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. Pearson accuracy correlation translates into Spearman ranking correlation and relative model order stability. Similarly, the overly easy gpt-4.1-mini synthetic benchmark in Figure 12 produces a mis-matched rank ordering (Spearman correlation $0.564$) in Figure 13(b). While ensembling various question generating LMs together will partially mitigate this effect, it is worth validating the set of generating models against the datasets in this paper to confirm the ensemble is not negatively impacting overall synthetic benchmark quality. The tendency of certain generating models to produce overly long, detailed questions that LMs find easier to answer may be the cause of the observed benchmark over-performance.
5.7 Multiple-choice vs open-ended questions
During this study, both multiple choice and open-ended questions were evaluated. None of our findings change whether using multiple choice questions for simplified evaluation benchmark grading, or open-ended questions to produce harder evaluation benchmarks where knowledge synthesis is required. The models that are good at creating synthetic multiple choice questions are also good at creating open ended questions. The models which create overly easy questions by including too much detail (like phi-4) do so for both open ended and MCQ, with the long question problem becoming even more pronounced for open ended questions as shown in Figure 9 in Appendix D.
Table 5 presents Pearson benchmark evaluation accuracy correlation and Spearman rank order correlation between the human and synthetic evaluations. See Appendix E for an expanded table showing correlations across all three LMs used for human question reformatting. Llama-4-Maverick was unexpectedly dominant in these final Table 5 results, even outperforming the ensemble of models (though not significantly).
Table 3: Pearson accuracy and Spearman rank correlation between synthetic and human benchmarks
Metric Pearson Accuracy Correlation Spearman Rank Correlation
Generating Model Multiple Choice Open Ended Multiple Choice Open Ended
Ensemble 0.7452 0.7095 0.9679 0.9209
Llama-3.2-3B-Instruct 0.4857 0.4709 0.6143 0.7143
Llama-3.3-70B-Instruct 0.6615 0.6034 0.9571 0.8813
Llama-3.1-8B-Instruct 0.6608 0.6217 0.7964 0.7275
Llama-4-Maverick 0.7891 0.7233 0.9679 0.9648
gemma-3-12b-it 0.6879 0.4085 0.8929 0.8198
gemma-3-27b-it 0.7477 0.6962 0.9571 0.9516
gemma-3-4b-it 0.6575 0.5787 0.6393 0.9209
gpt-4.1 0.6188 0.5216 0.9179 0.8549
gpt-4.1-mini 0.4015 0.5780 0.5643 0.8418
o4-mini 0.6542 0.5700 0.9250 0.7934
phi-4 0.1547 0.4965 0.6964 0.1780
Qwen2.5-7B-Instruct 0.5813 0.5905 0.7321 0.8901
5.8 Case study on two academic papers
To demonstrate our approach in a more realistic setting, we selected two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” [44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” [45]. The documents were chunked into 27 and 28 contexts respectively. “Current Solutions” averaged $7.7$ topics per context for a total of 210 questions. “Recent Advances” averaged $6.9$ topics per context for a total of 194 questions. For MCQ, the two synthetic benchmarks’ Spearman rank correlation is $0.911$, for OE the rank correlation is $0.969$. Under MCQ the gemma-3 models perform middle of the pack on these two benchmarks, but under OE questions the 12b and 27b outperform all other models. See Figure 14 in Appendix G for full rank ordering and examples of the generated questions.
5.9 Limitations
This work relies on a simple and possibly non-optimal document chunking approach. Extensions which cluster related topics or provide more fine-grained control over the information present in each chunk could improve the granularity and specificity of the generated benchmark. Including information found in tables, charts, and images would provide richer information sources that benchmark creation may take advantage of. Additionally, improving the nuance and difficulty of the questions significantly improves LM capability measurement. This likely requires careful management of the context information provided to the generating LM. Finally, the datasets used in this study to validate correlation with human written questions are all NLP extractive question answering type datasets. This limits the quality of the synthetic benchmarks as the grounding documents are fairly simplistic, especially compared to the two academic papers used in Section 5.8. Only the sec_qa dataset, which is grounded in an open source computer security textbook, comes close to the topic diversity and complexity of the example academic papers. All other validation datasets rely upon more general trivia type knowledge.
6 Conclusion
We introduce a methodology for creating synthetic benchmarks to evaluate language model (LM) capabilities grounded in high-quality document corpora selected by practitioners. This ensures the generated benchmarks are directly relevant to their specific domains and interests. This approach extends general-purpose human curated LM benchmarking into specialized domains, allowing for the identification of strengths and weaknesses that may not be apparent in broader evaluations.
Our method supports automatic generation of both multiple-choice questions (optimizing for ease of grading) and open-ended questions (optimizing for diagnostic relevance), allowing users to choose based on their evaluation goals. Validation against six human-curated NLP dataset achieves a Spearman rank correlation of $0.96$ and a benchmark Pearson accuracy correlation of $0.79$ between synthetic and human curated questions, suggesting that our approach is consistent with human-generated question/answer benchmarks. Additionally, we identify a potential synthetic data generation failure mode: very long questions can artificially inflate model performance, likely by overspecifying information related to the answer. This methodology enables the creation of targeted, relevant benchmark evaluations of domain-specific Language Model capabilities, facilitating deeper exploration into the capabilities and limitations of Language Models deployed for complex knowledge work.
Appendix A Expanded related works
This approach draws inspiration from NP-completeness, where easy to verify solutions may be hard to generate. The LM under evaluation has a harder job than the LM that converts existing documents into appropriately formatted test questions with accompanying answers. The evaluation LM must produce the correct information from just the question; by contrast, the benchmark creation LM has the source material. In other words, given the grounding paragraph, a language model should be much more capable at converting that textual information into a multiple choice or short answer question than it would be at zero-shot answering the very same question. This task difficulty discrepancy provides hope that this approach will produce viable results long term. Asking an LM to generate diagnostically useful evaluation benchmarks drawn only from information stored in model weights is difficult, but conversion of information from a high quality source into test questions is attainable.
This work draws on the rich history of language model evaluation, where LMs are already being used to judge and score the results of human curated benchmarks. When scoring outputs of an LM producing a multiple choice answer, it is often useful to have another LM reason about and extract the specified [A,B,C,D] value, instead of hoping the model under evaluation only produced the exact match answer letter [3, 4, 46]. This process becomes even more important when evaluating short answers from LMs instead of multiple choice responses, or when scoring the responses and providing reasoning as to why an answer was correct or incorrect [5, 6, 7]. This idea of automating question answer generation has been approached before. The most similar contemporary work is YourBench by Shashidhar et al. [8], which leverages the convergent idea of grounding documents supporting LM benchmark generation, demonstrating the synthetic replication of MMLU. Additionally, AutoBencher by Li et al. explores leveraging synthetic data (supported by privileged information) for model evaluation with RL to optimize the generated evaluation for discovering gaps in LM capability [9]. The use of grounding documents or privileged information greatly improves synthetic data benchmark utility, as the evaluation methodology no longer relies on simply asking an LM for questions (and answers).
Question from context and answer
Early Question Answer Generation (QAG) approaches attempted to solve the question creation from a context + correct answer input. These approaches include improvements in selecting better LM samples [47], increasing sample diversity over distributional shifts [48], explorations of interrogative prompting against the context [49], and evaluating round trip information consistency [50]. These early methods sometimes fine-tune generation models to improve task performance, though largely that is only done with smaller models [51, 52]. Some work focused on fine-tuning specialty models to improve option generation for multiple-choice questions [52]. Other early work during the shift to general purpose LMs involved leveraging those language models directly as data generation sources. This took the form of prompting the model to write questions (without grounding context) [53, 54, 55, 56], or configuring an LM-as-evaluator/interrogator dynamically interacting with the LM under evaluation to explore capability in a reference-free manner [53]. These approaches attempt to address the dataset leakage problem where benchmarks, once published, become part of the training dataset for the next round of models. As the capability of LMs increases over time, setting up adversarial LM-evaluator vs LM-evaluatee systems is becoming more feasible [57]. Nevertheless, extending those systems such that the LM-evaluator has grounding documents to pull from should strengthen the performance.
Scaffolding question generation with grounding documents
With the rise of general purpose token-tumblers like ChatGPT, longer and more complex grounding documents become increasingly useful. There is no longer a need to constrain synthetic benchmark creation to a short paragraph context + correct answer input; the LM can operate on larger and more complex document chunks, with less pre-processing required to make the input context data useful. Contemporaneous work YourBench [8] and AutoBencher [9] can leverage lightly-filtered web-pages as the context, chunking the data into rough paragraphs for ingestion into the question generation and grounding system. The SecQA benchmark was created partly through synthetic data grounded in an open-source computer security textbook [10]. Similarly (though much earlier), the PubMedQA benchmark was created through human and synthetic data generation grounded in biomedical research paper abstracts [11]. Some research demonstrates that operating purely on document summaries produces inferior generated questions compared to leveraging the full document [12]. With modern long-context LMs, summarization is becoming less of a problem, as models can manipulate significant volumes of text within their attention context. Synthetic data question generation systems have seen rapid improvement. This is partly driven by capability improvements in the LMs powering synthetic data generation. However, there has also been exploration into improvements in prompting for diverse question generation [13] and breadth of topic generation for chatbot arena style questions [14]. Enhancing retrieval augmented generation (RAG) systems for domain specific applications has also used synthetic data question generation to compare baseline RAG to fine-tuning [15].
LM as feature extractors
There are many aspects of text that one might want to quantify, however imprecisely. However, devoting the required copious human annotation hours is just impractical. Current LMs end up being acceptable feature extractors, when provided detailed prompts. Chiang and Lee demonstrate that LMs can substitute for human evaluators in assessing text quality with strong correlation in the results [16]. Similarly, Han et al. leverage LM evaluators for judging quality and find that results correlate highly with human judgments [17]. The use of LMs to extract features from text is not to dissimilar from using the LM-as-a-Judge [18] pattern for large scale open ended capability evaluations [18, 19, 21]. The LM is capable of grading open ended evaluation benchmark questions for correctness; therefore determining features other than answer correctness from a text sample is viable. However, there are limits on what LM-as-a-Judge is capable of ground truthing, especially in fields like information retrieval [20, 21] where LMs under-perform on more complex relevance judgments. Relying on LMs to extract properties like answer correctness given sufficient scaffolding and context is a well-used pattern.
Multiple choice vs open ended
When constructing knowledge or fact based evaluation benchmarks, there are two broad categories of answers: multiple choice and open ended. Open-ended questions provide the model less information and require synthesis [22], but at the cost of needing another LM to evaluate the correctness of the results (LM-as-a-Judge [18][23]. Multiple choice question (MCQ) results can be parsed and graded more simply. Modern evaluation frameworks like Inspect-AI [26] enable shuffling answer options and contain default prompts for both open-ended and MCQ evaluations. With MCQ, having the answers visible to the LM can artificially inflate performance [24], even going as far as the model identifying the correct answers without seeing the question [25]. Both open-ended and multiple choice questions have weaknesses to question reformatting and phrasing where semantically identical questions produce different answers [27]. The simplest form of open-ended evaluation relies on a judge LM, though other frameworks have been proposed which attempt to improve the precision of open-ended evaluations [28, 29].
Appendix B Impact of question and answer length
While longer more detailed questions (both MC and OE) produce a model over-performance (see Figure 5), the same trend does not appear in answer length. Figure 7 demonstrates this via scatterplot showing synthetic benchmark accuracy over-performance. Only the gpt-4.1-mini at the very right of the plot is significantly biased above the $y=0$ expected trend line. The over-performance trend shown for questions in Figure 5(b) is absent.
Figure 7: Impact of synthetic MCQ answer options token length on evaluation accuracy over-performance. LMs under evaluation do not over-perform when answer options are longer in the same way long questions cause over-performance.
The model gpt-4.1-mini demonstrates the largest over-performance for both question length and answer length as shown in Figure 5(b). The following question answer pair was generated by gpt-4.1-mini (37 tokens) using a context from Squadv2, highlighting the difference in model verbosity in both question and answers:
1"question": "How does the widespread accessibility and overuse of antibiotics, particularly in livestock, mechanistically contribute to the accelerated development of bacterial resistance, and what global health implications does this relationship entail?",
2
3"choices": {
4 "A": "Overuse increases bacterial exposure to antibiotics, promoting selection for resistant strains, which leads to treatment failures and a global rise in antimicrobial-resistant infections.",
5 "B": "Limited antibiotic use in livestock reduces bacterial mutations, thereby decreasing resistance and minimizing global health risks.",
6 "C": "Antibiotic accessibility restricts bacterial gene exchange, preventing resistance development and protecting global health.",
7 "D": "Overuse of antibiotics in humans alone causes resistance, while use in livestock has no significant impact on bacterial evolution or public health."},
8
9"answer": "A"
Contrast with this question answer pair generated by Llama-4-Maverick-17B-128E-Instruct using the same context and topic. Maverick produces a significantly shorter question (22 tokens) and answers:
1"question": "What is the primary mechanism by which the widespread use of antibiotics in livestock contributes to the development of antibiotic-resistant bacteria in humans?",
2
3"choices": {
4 "A": "By directly transferring resistant bacteria from animals to humans through the food chain.",
5 "B": "By exerting selective pressure that favors the survival and proliferation of resistant bacterial strains.",
6 "C": "By altering the human gut microbiota, thereby reducing its ability to fight off infections.",
7 "D": "By inducing genetic mutations in human pathogens that confer resistance."},
8
9"answer": "B"
Appendix C Synthetic benchmark question diversity
Most LMs produce similar diversity scores that are closely aligned with the human questions Some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 shows the question diversity per generating model along with the original question diversity.
Figure 8: Synthetic question diversity (pairwise cosine distance) per generating model. Violin bars show the spread across the 6 NLP datasets.
Appendix D Multiple-choice vs open-ended questions
The generating models which create overly easy questions by including too much detail (like phi-4 and gpt-4.1-mini) do so for both open ended and multiple choice questions. The long question problem becomes even more pronounced for open ended questions as shown in Figure 9.
(a) MC question token count by generating model.
(b) OE question token count by generating model.
Figure 9: Most generating models are within $20\%$ of human question token counts, but some produce overly long questions. The opened-ended questions trend toward longer questions overall.
[…]
(a) Impact of multiple-choice question token count on evaluation accuracy over-performance, caused by overly easy questions.
(b) Impact of open-ended question token count on evaluation accuracy over-performance, caused by overly easy questions.
Figure 10: Longer synthetic open-ended questions have the same impact on evaluation accuracy over-performance as MCQ.
While similar trends in question length and the over-performance that induces, the evaluation accuracy correlations and model rank order work equally well for open ended and multiple choice questions. Figure 11(a) plots the correlation in evaluation accuracy between the human reformatted questions and the synthetic generated benchmark for multiple choice, and Figure 11(b) for open-ended. Figure 11(a) shows a sec_qa($\diamond$) dataset specific oddity: the over-performance on human questions (top right) stems from the multiple choice human questions being very easy and produces an approximately flat line over various synthetic versions generated by LMs. Interestingly enough, this over-performance disappears under the open-ended question regime for this dataset as seen in Figure 11(b).
(a) Multiple Choice evaluation accuracy for the generating model ensemble.
(b) Open Ended evaluation accuracy for the generating model ensemble.
Figure 11: Multiple Choice vs Open Ended synthetic vs human benchmark evaluation performance exploring Pearson accuracy correlation. Both results demonstrate reasonable Pearson correlations between the benchmark evaluation accuracy.
Appendix E Evaluation benchmark accuracy and ranking correlation
Figure 11(a) shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the 5 bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of $0.745$ between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near $100\%$ evaluation accuracy (Pearson accuracy correlation $0.402$.). This over-performance is likely linked to the question length problem.
Figure 12: Synthetic over-performance for generating model gpt-4.1-mini indicating overly easy questions (Pearson accuracy correlation $0.402$.)
Figure 13 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. This matches the use case of a user selecting between various models for a specific application relevant to the provided grounding documents. Figure 13(a) is the rank order barchart equivalent of Figure 11(a) and Figure 13(b) is the rank order barchart equivalent of the failure case in Figure 12.
(a) Benchmark evaluation accuracy generated by Llama-4-Maverick-17B-128E-Instruct producing a Spearman rank correlation $0.968$.
(b) Benchmark evaluation accuracy generated by gpt-4.1-mini demonstrating over-estimation of LM capability with a Spearman rank correlation $0.564$.
Figure 13: Comparison of benchmark rank ordering between synthetic and human evaluation accuracy.
Table 4 showcases the Spearman rank correlation between evaluation benchmarks written by humans and the synthetic data versions. The table includes data for all three of the human question reformatting models used to demonstrate the minimal impact that reformatting has on the overall high level results. Overall, the best outcome is to ensemble together a few questions generated by various reasonably capable LMs to avoid the known slight self-bias/self-preference issue. While we used Llama-4-Maverick as the text feature extractor for measures like groundedness and clarity, that model was unexpectedly dominant in these final results.
Table 4: Spearman rank correlation between synthetic and human benchmarks for multiple choice (MC) and open-ended (OE) evaluations across generating models. Ensemble = bolded models.
Reformat Model Llama-4-Maverick gemma-3-27b-it o4-mini
Generating Model MC Rank OE Rank MC Rank OE Rank MC Rank OE Rank
Ensemble 0.9679 0.9209 0.9429 0.9253 0.9821 0.8813
Llama-3.2-3B-Instruct 0.6143 0.7143 0.6929 0.7143 0.6536 0.6791
Llama-3.3-70B-Instruct 0.9571 0.8813 0.9643 0.8681 0.9714 0.8242
Llama-3.1-8B-Instruct 0.7964 0.7275 0.7321 0.8110 0.7714 0.7275
Llama-4-Maverick 0.9679 0.9648 0.9536 0.9516 0.9821 0.9209
gemma-3-12b-it 0.8929 0.8198 0.8964 0.8418 0.9000 0.7890
gemma-3-27b-it 0.9571 0.9516 0.9857 0.9473 0.9714 0.9121
gemma-3-4b-it 0.6393 0.9209 0.6571 0.8989 0.6857 0.8637
gpt-4.1 0.9179 0.8549 0.9250 0.8725 0.9357 0.8110
gpt-4.1-mini 0.5643 0.8418 0.6286 0.8549 0.6071 0.7934
o4-mini 0.9250 0.7934 0.9036 0.8374 0.9429 0.7714
phi-4 0.6964 0.1780 0.6250 0.3011 0.7607 0.1736
Qwen2.5-7B-Instruct 0.7321 0.8901 0.7143 0.8945 0.7786 0.8505
Table 5 contains the benchmark evaluation Pearson accuracy correlation between the reformatted human questions and the synthetic data generated questions. Similarly to the Spearman ranking, Llama-4-Maverick performed very well in these tests. However, unlike the ranking data, using a ensemble of models produces significantly worse Pearson accuracy correlation numbers under some circumstances. Likely due to the long and overly easy questions written by gpt-4.1 and o4-mini.
Table 5: Pearson accuracy correlation between synthetic and human benchmarks for multiple choice (MC) and open-ended (OE) evaluations across generating models. Ensemble = bolded models.
Reformat Model Llama-4-Maverick gemma-3-27b-it o4-mini
Generating Model MC Corr OE Corr MC Corr OE Corr MC Corr OE Corr
Ensemble 0.7452 0.7095 0.7248 0.5630 0.6569 0.4798
Llama-3.2-3B-Instruct 0.4857 0.4709 0.5116 0.6181 0.4448 0.5083
Llama-3.3-70B-Instruct 0.6615 0.6034 0.7284 0.7464 0.6390 0.5798
Llama-3.1-8B-Instruct 0.6608 0.6217 0.7231 0.6776 0.6472 0.5795
Llama-4-Maverick 0.7891 0.7233 0.8755 0.8217 0.8109 0.7833
gemma-3-12b-it 0.6879 0.4085 0.7748 0.5809 0.6986 0.4617
gemma-3-27b-it 0.7477 0.6962 0.8499 0.8292 0.7426 0.7046
gemma-3-4b-it 0.6575 0.5787 0.7117 0.6932 0.6524 0.6040
gpt-4.1 0.6188 0.5216 0.7218 0.6760 0.6528 0.5236
gpt-4.1-mini 0.4015 0.5780 0.4795 0.6508 0.4190 0.5271
o4-mini 0.6542 0.5700 0.7346 0.6430 0.6904 0.5503
phi-4 0.1547 0.4965 0.2348 0.4951 0.2379 0.4043
Qwen2.5-7B-Instruct 0.5813 0.5905 0.6448 0.7199 0.5062 0.6019
Appendix F Question clarity and groundedness
We observed two failure modes with respect to grounding: Either the model hallucinates content absent from and not relevant to the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). For example, phi-4 generated the following question from the provided context:
1"question": "How did the educational policies of the British Empire during the late 19th and early 20th centuries shape the leadership qualities necessary for navigating modern challenges in the 21st century?"
2
3"context": "Victoria later described her childhood as \"rather melancholy\". Her mother was extremely protective, and Victoria was raised largely isolated from other children under the so-called \"Kensington System\", an elaborate set of rules and protocols devised by the Duchess and her ambitious and domineering comptroller, Sir John Conroy, who was rumoured to be the Duchesss lover. The system prevented the princess from meeting people whom her mother and Conroy deemed undesirable (including most of her fathers family), and was designed to render her weak and dependent upon them. The Duchess avoided the court because she was scandalised by the presence of King Williams bastard children, and perhaps prompted the emergence of Victorian morality by insisting that her daughter avoid any appearance of sexual impropriety. Victoria shared a bedroom with her mother every night, studied with private tutors to a regular timetable, and spent her play-hours with her dolls and her King Charles spaniel, Dash. Her lessons included French, German, Italian, and Latin, but she spoke only English at home."
Appendix G Case study on two academic papers
Our approach was demonstrated on two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” (labeled annurev-control-071020-104336[44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” (labeled arXiv_2502_17521v1[45]. Figure 14 showcases LM benchmark evaluation performance on these two datasets; the MCQ results show most models performing well before a sharp dropoff with 4B and fewer parameter models and the Open-Ended results with a much smoother decline in performance. Multiple-choice questions can be easier to answer due to the presence of the answer options, even if the evaluation framework shuffles the option order. This effect is likely the cause of the MCQ evaluation accuracy to remaining higher for longer.
(a) Multiple Choice Synthetic Questions
(b) Open-Ended Synthetic Questions
Figure 14: Evaluation benchmark rank ordering comparing MCQ to Open-Ended questions for two papers drawn from current research work [44, 45].
The following example questions were drawn from the MCQ synthetic benchmark constructed from arXiv_2502_17521v1 [45].
1{"question": "What is the primary distinction between the \"temporal cutoff\" approach and the \"rule-based generation\" approach in dynamic benchmarking?",
2
3"choices": {
4 "A": "Temporal cutoff relies on newly released information, while rule-based generation creates novel evaluation data points using predefined rules.",
5 "B": "Temporal cutoff uses predefined rules, while rule-based generation relies on newly released information.",
6 "C": "Temporal cutoff is a hybrid approach combining multiple methodologies, while rule-based generation is a standalone method.",
7 "D": "Temporal cutoff generates data points using LLMs, while rule-based generation uses manual curation."},
8
9"answer": "A"},
10
11{"question": "What is a potential future research direction for improving the design and standardization of dynamic benchmarking methods for Large Language Models (LLMs), given the current limitations and proposed evaluation criteria?",
12
13"choices": {
14 "A": "Developing more sophisticated data regeneration techniques to further minimize data contamination.",
15 "B": "Enhancing static benchmarking methods with more robust data encryption.",
16 "C": "Proposing new model architectures to reduce reliance on Internet-sourced training data.",
17 "D": "Focusing solely on post-hoc contamination detection methods."},
18
19"answer": "A"}
The following example questions were drawn from the Open-Ended synthetic benchmark created for arXiv_2502_17521v1.
1{"question": "How might the presence of contaminated benchmarks influence the trajectory of Large Language Model research and development, particularly in terms of model comparisons and deployment decisions?",
2
3"answer": "By leading to misleading conclusions that favor contaminated models."},
4
5{"question": "What are the implications of using LLMs to rewrite benchmark samples while preserving their original difficulty levels, as seen in methodologies like ITD, on the overall diversity and complexity of the resulting benchmark datasets?",
6
7"answer": "The resulting datasets may maintain complexity but risk lacking diversity if not combined with other rewriting strategies."}